### v1

# GS-SDF: LiDAR-Augmented Gaussian Splatting and Neural SDF for Geometrically Consistent Rendering and Reconstruction

Abstract
Digital twins are fundamental to the development of autonomous driving and embodied artificial intelligence. However, achieving high-granularity surface reconstruction and high-fidelity rendering remains a challenge. Gaussian splatting offers efficient photorealistic rendering but struggles with geometric inconsistencies due to fragmented primitives and sparse observational data in robotics applications. Existing regularization methods, which rely on render-derived constraints, often fail in complex environments. Moreover, effectively integrating sparse LiDAR data with Gaussian splatting remains challenging. We propose a unified LiDAR-visual system that synergizes Gaussian splatting with a neural signed distance field. The accurate LiDAR point clouds enable a trained neural signed distance field to offer a manifold geometry field, This motivates us to offer an SDF-based Gaussian initialization for physically grounded primitive placement and a comprehensive geometric regularization for geometrically consistent rendering and reconstruction. Experiments demonstrate superior reconstruction accuracy and rendering quality across diverse trajectories. To benefit the community, the codes will be released at https://github.com/hku-mars/GS-SDF.
I Introduction
(a) Data Collection Trajectory and Input Point Clouds
(b) Surface Reconstruction
(c) Geometrically Consistent Gaussian Splats
(d) Novel View Synthesis
Figure 1: We show the input raw colorized point clouds (FAST-LIVO2 SYSU Scene) collected by a real-world LiDAR-visual sensor system. We demonstrate the proposed system, GS-SDF, with the surface reconstruction results (color indicates the normal direction), Gaussian splats’ geometric distribution, and novel view synthesis results.
3D surface reconstructions and photorealistic renderings are essential for a wide range of applications, like augmented reality[1], and embodied artificial intelligence[2]. The increasing accessibility of cameras and cost-effective LiDAR sensors on robots has enabled the collection of rich multimodal data. LiDAR-visual SLAM technologies[3, 4, 5] are widely applied in 3D reconstruction tasks. However, these methods typically produce only discrete colorized raw point clouds, which lacks complete surface structure and rendering ability for view-dependent novel view synthesis. In practical digital-twin applications [6], high-quality watertight surface reconstructions and photorealistic rendering are critical.
Neural Radiance Fields (NeRFs) [7] have gained attention for their ability to generate high-fidelity renderings. However, their computational intensity and inefficiency in real-time applications limit their practicality. 3D Gaussian Splatting (3DGS) [8] has emerged as a compelling alternative to NeRFs for efficient, high-quality rendering. Unlike NeRF requiring laborious volume rendering, 3DGS employs explicit ellipsoidal Gaussian primitives for efficient rasterization. However, the fragmented structure of these primitives compromises structural continuity and often misaligns with underlying geometry. This misalignment hinders geometrically consistent rendering, particularly in robotics applications where free-view trajectories yield insufficient observations. Furthermore, under image-only supervision with imperfect camera poses, 3DGS frequently overfits to photometric cues, producing floating artifacts. It motivates us to introduce geometric regularization into the 3DGS framework.
Geometric inconsistencies in 3DGS manifest as rendering distortions. Current regularization strategies focus on rendering-depth-derived constraints, such as rendering normal consistency from 2D Gaussian splatting (2DGS) [9] or multi-view alignment from PGSR[10]. They focus on object-centric scenes with abundant multi-view regularization and limited generalization to complex environments. Depth cameras and LiDARs providing direct structural priors inspire us to incorporate these sensors into the 3DGS framework for geometric regularization. And depth cameras with limited precision are confined for indoor applications[11]. While LiDAR providing sparse point clouds remains challenging to integrate with NeRF[12] or 3DGS[13] frameworks designed for dense, image-scale geometry.
The combination of LiDAR and neural signed distance fields (NSDF) has shown promise in high-granularity surface reconstructions[14]. The NSDF, providing a manifold geometry field, motivates us to explore a more comprehensive geometric regularization for Gaussian splats and the importance of the Gaussian initialization to the results. While, surface reconstruction from Gaussian splats is complicated by their discontinuous nature. The truncated Signed Distance Field (TSDF) fusion [15] is commonly applied for 3DGS-based surface reconstructions through depth images rendered from Gaussian splatting, which demonstrates unsatisfactory inaccuracies in surface reconstruction for generalized scenes.
To address above limitations, we propose the combination of 2D Gaussian splatting and neural signed distance field for LiDAR-visual systems. We aim to reconstruct both the appearance and surface of generalized scenes using posed images and low-cost LiDAR data from any casual trajectories. We emphasize the importance role of the Gaussian initialization in returning a good structure of the Gaussian splats. We explore the combination of the NSDF and Gaussian splats to leverage the structural and photometric hints for high-granularity surface reconstruction and geometrically consistent rendering, as shown in Fig. 1. Our contributions are as follows:
  1. 1.
    A unified LiDAR-visual system achieving geometrically consistent photorealistic rendering and high-granularity surface reconstruction.
  2. 2.
    A physically grounded Gaussian initialization based on the neural signed distance field to boost training convergency and reduce floating artifacts.
  3. 3.
    A thorough shape regularization gives bidirectional supervision between the neural signed distance field and Gaussian splats for geometrically consistent rendering and reconstruction.
Extensive experiments validate our approach across diverse scenarios, demonstrating superior reconstruction accuracy and rendering quality.
II Related Works
II-A Geometrically Consistent Novel View Synthesis
NeRFs [7] and 3DGS [8] have been widely employed for novel view synthesis, owing to their remarkable photorealistic rendering capabilities. However, neither of them can guarantee geometric consistency, frequently leading to the emergence of floating artifacts [16]. To tackle this problem, geometric regularizations and structural priors derived from depth sensors have been incorporated into both NeRF and 3DGS. 2DGS [9] projects Gaussian ellipsoids into surfels and enforces normal-depth alignment geometric regularization. PGSR [10] provides further guarantees for multi-view rendering consistency regularization. RGBD cameras that provide aligned color and depth information are well practiced in both NeRFs [17, 18] and 3DGS [11] to align rendered depth images with sensor-acquired ones. However the low measurement accuracy of depth cameras limited their application in indoor scene settings. LiDAR sensors provide accurate depth information, but their sparse nature complicates integration with NeRFs [19] or 3DGS [20] frameworks designed for dense, image-scale geometry. M2Mapping [12] enforces geometric consistency to NeRF through the combination of NSDF and ray marching, albeit at a significant computational expense. LIV-GaussMap [20] stabilizes splat orientations using planar priors derived from LiDAR-inertial odometry. And LI-GS [13] utilizes Gaussian Mixture Models for constraining the Gaussian splats’ distribution.
II-B Surface Reconstruction
Surface reconstructions are predominated using point clouds with direct structural information. The TSDF fusion [15] is well practiced for surface reconstructions[21]. While the quest for continuous surface reconstruction motivates the exploration of implicit representations, like Poisson functions [22] or NSDFs [23, 24]. In terms of the exploration of Gaussian splats for surface reconstruction, its discontinuous nature complicates surface reconstruction. Recent approaches circumvent this limitation through intermediate volumetric representations. SuGaR [25] computes an approximate density field based on the opacities of neighboring Gaussians, and GOF [26] defines per-point opacity as the minimal opacity across training views. However, both density and opacity fields suffer from computational inefficiency and ambiguous surface level sets for surface reconstruction. In contrast, the SDF facilitates precise surface extraction through defined zero-level sets. The TSDF fusion driven by the rendered depth images is favored by 3DGS-based surface reconstructions[9, 10], but yields insufficient accuracy. Image-driven NSDFs combined with 3DGS [27, 28] are still under exploration for achieving considerable accuracy in surface reconstructions.
III Methodology
Figure 2: The overall pipeline of the proposed system, GS-SDF, for geometrically consistent rendering and reconstruction.
Given posed images and point clouds from LiDAR-inertial-visual odometry (e.g., FAST-LIVO2 [5]), we propose a framework to jointly reconstruct high-granularity scene geometry and photorealistic appearance. As shown in Fig. 2, our pipeline consists of three stages: A neural signed distance field (NSDF) is first trained using point clouds to establish a manifold geometry field (Sec. III-B). Gaussian splats are initialized from the NSDF (Sec. III-C) and a thorough SDF-aided shape regularization for both Gaussian splats and the NSDF (Sec. III-D) achieve a geometric-consistent rendering and reconstruction.
III-A 2D Gaussian Splatting
2DGS [9] represents 3D scenes using planar Gaussian disks. Each disk is characterized by a center $\boldsymbol{p}\in\mathbb{R}^{3}$, orthogonal tangent vectors $\{\boldsymbol{t}_{u},\boldsymbol{t}_{v}\in\mathbb{R}^{3}\}$ that define its local plane, scale factors $\{s_{u},s_{v}\in\mathbb{R}\}$, opacity $\alpha\in\mathbb{R}$, and view-dependent color $\boldsymbol{c}\in\mathbb{R}^{3}$ encoded using spherical harmonics. The disk’s normal vector is computed as the cross product of its tangent vectors: $\boldsymbol{n}=\boldsymbol{t}_{u}\times\boldsymbol{t}_{v}$. A point $\boldsymbol{u}=[u,v]^{\top}$ in the tangent plane is mapped to 3D space via the transformation $\boldsymbol{p}(u,v)=\boldsymbol{p}+s_{u}u\boldsymbol{t}_{u}+s_{v}v\boldsymbol{t}_{v}$, with the Gaussian weight kernel defined as $\mathcal{G}(\boldsymbol{u})=\exp\left(-\frac{u^{2}+v^{2}}{2}\right)$. For rendering, Gaussians are transformed into camera coordinates and depth-sorted. The final pixel color $\boldsymbol{c}$ is computed through alpha blending:
$\boldsymbol{C}=\sum_{i=1}\boldsymbol{c}_{i}\alpha_{i}\mathcal{G}_{i}(\boldsymbol{u})\prod_{j=1}^{i-1}\left(1-\alpha_{j}\mathcal{G}_{j}(\boldsymbol{u})\right)=\sum_{i=1}\boldsymbol{c}_{i}w_{i}.$ (1)
The Gaussian attributes are optimized using a combined rendering loss function of L1 loss and structural dissimilarity (DSSIM) metrics [8]:
$\mathcal{L}_{c}=0.8L_{1}({\boldsymbol{C}},\bar{\boldsymbol{C}})+0.2L_{\mathrm{DSSIM}}({\boldsymbol{C}},\bar{\boldsymbol{C}}),$ (2)
where $\bar{\boldsymbol{C}}$ denotes the ground truth pixel color.
III-B Neural Signed Distance Field
We utilize hash encoding [29] and multi-layer perceptrons (MLPs) to construct a neural network for shaping a scalable signed distance field: $({s},{\beta})=f(\boldsymbol{x})$. A 3D point $\boldsymbol{x}\in\mathbb{R}^{3}$ is mapped to a signed distance value $s\in\mathbb{R}$ and a scale factor $\beta\in\mathbb{R}$[12]. For a LiDAR measurement comprising an origin ${}^{L}\boldsymbol{o}$, a ray direction ${}^{L}\boldsymbol{d}$ and a distance $t$, we uniformly sample points along the ray path $\{t_{i}:{}^{L}\boldsymbol{x}_{i}={}^{L}\boldsymbol{o}+t_{i}\ {}^{L}\boldsymbol{d}\}$. The NSDF gives prediction at each sampled point: $(s_{i},\beta_{i})=f({}^{L}\boldsymbol{x}_{i})$. The ray distance $\bar{s}_{i}=t-t_{i}$ derived from the sampled distance is converted to an occupancy value $\bar{o}_{i}=\Phi(-\bar{s}_{i},\beta_{i})$ using the sigmoid function $\Phi(v,h)=(1+\exp(-v/h))^{-1}$ and the predicted scale factor ${\beta}_{i}$. The NSDF is trained using binary cross-entropy loss [24]:
$\mathcal{L}_{sdf}=-\sum_{i=1}\big{[}{o}_{i}\log(\bar{o}_{i})+(1-{o}_{i})\log(1-\bar{o}_{i})\big{]},$ (3)
where $o_{i}=\Phi(-{s}_{i},{\beta}_{i})$ is the predicted occupancy value. The Eikonal regularization [30] is empolyed to ensure the gradient of the SDF maintains unit magnitude:
$\mathcal{L}_{e}=\sum_{i=1}\left(\|\nabla f(\boldsymbol{x}_{i})\|_{2}-1\right)^{2}.$ (4)
III-C Initialization of Gaussian Splats
In this section, we address the crucial role of the Gaussian initialization plays a in optimizations and propose a careful Gaussian initialization with the help of a coarse neural signed distance field trained using input point cloud. We provide a thorough initialization of Gaussians boosts the convergency efficiency and suppresses floating artifacts.
III-C1 SDF-aided Geometry Initialization
The neural signed distance field, providing a manifold geometry field, can well stabilize the initial geometry of Gaussians. We extract a surface mesh from the NSDF using marching cubes[31], and utilize its vertices as initial Gaussian positions $\boldsymbol{p}$. This approach provides spatially accurate and denser positional initialization compared to structure-from-motion points, and reduces the impact of sensor outliers compared to the sampling initialization from input point clouds. The marching cube resolution directly determines the initial scale $\{s_{u},s_{v}\}$ of the Gaussian splats. We leverage the SDF’s geometric hints and Gram-Schmidt orthogonalization for Gaussian orientation initialization, using gradient direction $\boldsymbol{n}=\frac{\nabla f}{\|\nabla f\|}$ as Gaussian normals and principal curvature direction $\boldsymbol{b}=\frac{\nabla^{2}f}{\|\nabla^{2}f\|}$ to define the second rotation axis $\boldsymbol{t}_{u}=\frac{\boldsymbol{b}-(\boldsymbol{n}^{T}\boldsymbol{b})\boldsymbol{n}}{\|\boldsymbol{b}-(\boldsymbol{n}^{T}\boldsymbol{b})\boldsymbol{n}\|}$, and the first rotation axis $\boldsymbol{t}_{v}=\boldsymbol{n}\times\boldsymbol{t}_{u}$ for the final rotation: $\boldsymbol{R}=\left[\boldsymbol{t}_{u},\boldsymbol{t}_{v},\boldsymbol{n}\right]$. The opacity is derived from the SDF value and Gaussian kernel: $\alpha=\exp(-\frac{s^{2}}{\beta})$.
III-C2 Sky Initialization
For the infinite background, we uniformly tile the opaque Gaussian over a map-sized sphere. It provides a canvas for infinite objects to avoid floating artifacts derived from the foreground Gaussians.
III-C3 Color Initialization
Next, we maintain fixed Gaussians’ structural attributes (position, rotation, scale, and opacity) while training the Gaussians on all training images in a single round for color initialization. The color initialization helps to solid the Gaussians’ structural from being deviated in early training stages.
III-D Geometric Regularization for Rendering
The rasterization-based rendering of Gausisan splatting can lead to occlusion artifacts where foreground elements inappropriately obscure background content. Geometric regularization are designed to hold the rendering consistency of different views to avoid such artifacts. Considering that the LiDAR structure priors are not always available everywhere, we incorporate both the rendering-based regularization in 2D space and the SDF-aided shape regularization in 3D space to guarantee a geometrically consistent rendering.
III-D1 Rendering Consistency Regularization
Following with the rendering regularization in 2DGS[9], we obtain the rendered normal images from splats’ normals: $\boldsymbol{N}=\sum_{i=1}\boldsymbol{n}_{i}w_{i}$. And the finite differential normal images from the rendered depth images: $\hat{\boldsymbol{N}}(x,y)=\frac{\nabla_{x}\boldsymbol{p}_{d}\times\nabla_{y}\boldsymbol{p}_{d}}{\|\nabla_{x}\boldsymbol{p}_{d}\times\nabla_{y}\boldsymbol{p}_{d}\|}$, where $\boldsymbol{p}_{d}$ is the point derived from the rendered depth images at piexl $(x,y)$. The rendering consistency regularization aligns the rendered normal with the finite differential normal as follows:
$\mathcal{L}_{r}=(1-\boldsymbol{N}^{T}\hat{\boldsymbol{N}}).$ (5)
III-D2 SDF-aided Shape Regularization
With the NSDF providing a manifold geometry field, we can regularize the Gaussian splats’ shape to align with the physical surface. Consider that the splats represent a local plane of the surface, simply pulling the splat center $\boldsymbol{p}$ to the zero-level sets is insufficient[28], as shown in Fig. 3(a). We propose the shape regularization to enables comprehensive structural regularization across the entire splat surface. For each splat, we sample a point $\boldsymbol{u}_{s}=[u_{s},v_{s}]^{T}\sim\mathcal{N}(0,1)$ on its disk surface:
$\boldsymbol{p}_{s}$ $=\boldsymbol{p}+\boldsymbol{t}_{u}s_{u}u_{s}+\boldsymbol{t}_{v}s_{v}v_{s}.$ (6)
For $i$-th sampled point $\boldsymbol{p}_{s_{i}}$, we infer its signed distance value $f(\boldsymbol{p}_{s_{i}})$ to align splats with NSDF’s zero-level sets:
$\mathcal{L}_{s}=\sum_{i=1}\frac{1}{2}W_{i}\mathcal{G}(\boldsymbol{u}_{s_{i}})f(\boldsymbol{p}_{s_{i}})^{2},$ (7)
where $W_{i}=\sum_{k}w_{k}$ represents the accumulated weight of the $i$-th splat from Eq. 1. Considering that the SDF gradient maintains unit magnitude $\|\frac{\partial f(\boldsymbol{p}_{s_{i}})}{\partial\boldsymbol{p}_{s_{i}}}\|=1$, we adopt the $L_{2}$ loss function to enable adaptive refinement with proper gradients: $\frac{\partial\mathcal{L}_{s}}{\partial\boldsymbol{p}_{s_{i}}}=W_{i}\mathcal{G}(\boldsymbol{u}_{s_{i}})f(\boldsymbol{p}_{s_{i}})\frac{\partial f(\boldsymbol{p}_{s_{i}})}{\partial\boldsymbol{p}_{s_{i}}}$.
(a) Center Regularization
(b) Shape Regularization
Figure 3: Illustration of different geometric regularization. The blue splats represent the initial splat, and the orange splats represent the optimized splat. The dashed lines represent the optimization directions.
As illustrated in Fig. 3(b), we take one sampling for each Gaussian splat in the shape regularization. This regularization provides a concise all-in-one shape alignment to enable gradient-based supervision of all structural attributes through the following partial derivatives: $\frac{\partial\boldsymbol{p}_{s}}{\partial\boldsymbol{p}}=\boldsymbol{I}$, $\frac{\partial\boldsymbol{p}_{s}}{\partial\boldsymbol{t}_{u}}=s_{u}u_{s}$, $\frac{\partial\boldsymbol{p}_{s}}{\partial\boldsymbol{t}_{v}}=s_{v}v_{s}$, $\frac{\partial\boldsymbol{p}_{s}}{s_{u}}=\boldsymbol{t}_{u}u_{s}$, $\frac{\partial\boldsymbol{p}_{s}}{s_{v}}=\boldsymbol{t}_{v}u_{s}$.
III-E Optimization
The overall optimization function is a combination of the supervisions from point clouds, images, and geometric regularizations:
$\mathcal{L}$ $=\mathcal{L}_{sdf}+\lambda_{e}\mathcal{L}_{e}+\mathcal{L}_{c}+\lambda_{r}\mathcal{L}_{r}+\lambda_{s}\mathcal{L}_{s},$ (8)
where $\lambda_{e}=0.1$, $\lambda_{r}=0.01$, and $\lambda_{s}=0.005$ are hyperparameters that balance the loss terms.
TABLE I: Quantitative surface reconstruction, interpolation(I) and extrapolation(E) rendering results on the Replica dataset.
Metrics Methods Office-0 Office-1 Office-2 Office-3 Office-4 Room-0 Room-1 Room-2 Avg.
C-L1[cm]$\downarrow$ VDBFusion 0.618 0.595 0.627 0.685 0.633 0.637 0.558 0.656 0.626
iSDF 2.286 4.032 3.144 2.352 2.120 1.760 1.712 2.604 2.501
SHINE-Mapping 0.753 0.663 0.851 0.965 0.770 0.650 0.844 0.826 0.790
H2Mapping 0.557 0.529 0.585 0.644 0.616 0.568 0.523 0.616 0.580
M2Mapping 0.494 0.476 0.501 0.517 0.531 0.486 0.455 0.532 0.499
Ours 0.500 0.487 0.502 0.525 0.540 0.489 0.460 0.542 0.506
F-Score[%$\uparrow$ VDBFusion 97.129 97.107 97.007 96.221 96.993 97.871 98.311 96.117 97.095
iSDF 75.829 55.429 71.232 77.274 81.281 78.253 84.801 76.815 75.114
SHINE-Mapping 94.397 95.606 92.150 87.855 94.427 97.010 91.307 92.601 93.169
H2Mapping 98.430 98.279 97.935 97.077 97.391 98.850 99.003 96.904 97.983
M2Mapping 98.970 98.771 98.657 98.415 98.158 99.238 99.496 97.677 98.673
Ours 98.986 98.628 98.671 98.290 97.990 99.242 99.468 97.582 98.607
SSIM(I)$\uparrow$ InstantNGP 0.981 0.980 0.963 0.960 0.966 0.964 0.964 0.967 0.968
3DGS 0.987 0.980 0.980 0.976 0.980 0.977 0.982 0.980 0.980
2DGS 0.978 0.958 0.967 0.964 0.969 0.966 0.971 0.969 0.968
PGSR 0.981 0.967 0.966 0.963 0.966 0.898 0.973 0.972 0.961
H2Mapping 0.963 0.960 0.931 0.929 0.941 0.914 0.929 0.935 0.938
MonoGS 0.985 0.982 0.973 0.970 0.976 0.971 0.976 0.977 0.975
M2Mapping 0.985 0.983 0.973 0.970 0.977 0.968 0.975 0.977 0.976
Ours 0.986 0.971 0.979 0.976 0.980 0.978 0.980 0.979 0.979
PSNR(I)$\uparrow$ InstantNGP 43.538 44.340 38.273 37.603 39.772 37.926 38.859 39.568 39.984
3DGS 43.932 43.237 39.182 38.564 41.366 39.081 41.288 41.431 41.010
2DGS 41.854 40.796 38.138 37.704 40.144 37.484 39.905 39.748 39.472
PGSR 43.018 42.303 38.405 37.634 40.143 38.165 40.246 40.415 40.041
H2Mapping 38.307 38.705 32.748 33.021 34.308 31.660 33.466 32.809 34.378
MonoGS 43.648 43.690 37.695 37.539 40.224 37.779 39.563 40.134 40.034
M2Mapping 44.369 44.935 39.652 38.874 41.318 38.541 40.775 40.705 41.146
Ours 44.383 43.389 39.851 39.245 41.989 39.303 41.500 41.802 41.433
SSIM(E)$\uparrow$ InstantNGP 0.972 0.961 0.934 0.938 0.952 0.918 0.936 0.941 0.944
3DGS 0.936 0.897 0.924 0.917 0.925 0.881 0.915 0.919 0.914
2DGS 0.944 0.829 0.922 0.920 0.940 0.901 0.924 0.922 0.913
PGSR 0.949 0.870 0.932 0.931 0.938 0.970 0.924 0.930 0.930
H2Mapping 0.957 0.955 0.932 0.925 0.937 0.866 0.916 0.917 0.926
MonoGS 0.974 0.961 0.945 0.942 0.950 0.912 0.942 0.946 0.947
M2Mapping 0.980 0.976 0.960 0.964 0.970 0.955 0.963 0.965 0.967
Ours 0.973 0.952 0.956 0.952 0.958 0.945 0.949 0.954 0.955
PSNR(E)$\uparrow$ InstantNGP 39.874 39.120 31.274 32.135 34.458 32.587 33.024 32.266 34.341
3DGS 31.220 29.959 27.411 26.442 28.324 27.541 28.429 27.139 28.307
2DGS 30.847 25.546 27.248 25.499 29.270 28.149 28.694 26.846 27.762
PGSR 32.525 29.544 28.243 27.032 28.944 27.801 28.977 27.603 28.834
H2Mapping 36.740 37.841 31.427 31.144 31.988 28.815 31.192 30.603 32.468
MonoGS 39.197 38.818 29.740 29.664 31.632 29.949 31.126 30.621 32.593
M2Mapping 41.965 42.215 35.056 37.465 38.667 36.427 37.294 36.722 38.226
Ours 39.102 38.174 30.769 31.503 33.211 32.253 32.226 32.566 33.725
IV Experiments
IV-A Implementation Details
We represent our neural signed distance field following a similar architecture to InstantNGP [29], utilizing a combination of multi-resolution hash encoding and a tiny MLP decoder. Given any position $x$, the hash encoding concatenates each resolution’s interpolation features to form a feature vector of size $32$. A geometry MLP with 64-width and 3 hidden layers decodes the encoding feature to obtain the SDF value and scale. We sample 4096 rays for NSDF training with 4 surface samples and 4 free space samples on each ray. We first take 10000 iterations for NSDF training and leverage the trained NSDF for our initialization as described in Sec. III-C. Then, we follow the setting of 2DGS[9] to take 30000 iterations for the training. The implementation is based on LibTorch and CUDA, and all experiments are conducted on a platform equipped with an Intel i7-13700K CPU and an NVIDIA RTX 4090 GPU.
(a) VDB-Fusion
(b) iSDF
(c) SHINE-Mapping
(d) 2DGS
(e) PGSR
(f) H2Mapping
(g) M2Mapping
(h) Ours
(i) Ground Truth
Figure 4: Reconstructed mesh on the Replica dataset’s Office-2 and colors indicate the direction of the surface normal. Our method can capture more precise geometric details in slim objects.
(a) InstantNGP
(b) 3DGS
(c) 2DGS
(d) PGSR
(e) MonoGS
(f) H2Mapping
(g) M2Mapping
(h) Ours
(i) Ground Truth
Figure 5: Extrapolation rendering results on the Replica dataset’s room-0. Our method can capture more precise geometric details in slim objects and retain both the structure and texture in extrapolation rendering.
IV-B Experiments Settings
IV-B1 Baselines
We compare our method with other prevailing approaches. For pure surface reconstruction methods, we include the voxel-based method VDBFusion [21], and the neural-network-based method SHINE-Mapping [24]. For novel view synthesis methods, we include the NeRF-based method: InstantNGP [29], and the 3DGS-based methods initialized with the LiDAR point clouds: 3D Gaussian Splatting [8], 2D Gaussian Splatting [9], and PGSR[10]. For depth-aided methods, we include RGBD-based methods: H2Mapping [18] and MonoGS [11], and LiDAR-augmented NeRF-based method M2Mapping[12].
IV-B2 Metrics
To evaluate the quality of the geometry, we recover the NSDF to a triangular mesh using marching cubes [31] and calculate the Chamfer Distance (C-L1, cm) and F-Score ($<2$ cm, %) with the ground truth. For rendering quality evaluation, we use the Structural Similarity Index (SSIM) and Peak Signal-to-Noise Ratio (PSNR) to compare the rendered image with the ground truth image.
(a) Campus
(b) Sculture
(c) Culture
(d) Drive
(e) Station
(f) CBD
Figure 6: We show our surface reconstruction (color indicates the normal direction) and rendering results on the FAST-LIVO2 Datasets collected by real-world LiDAR-visual sensor systems under different trajectories, where from top to bottom are collected point clouds, reconstructed meshes, and novel view synthesis by Gaussian splatting.
IV-C Comparative Study
IV-C1 Replica Datasets
The Replica datasets [6] provides room-scale indoor simulation RGBD sensor data. We follow the M2Mapping[12] to emphasize the issues of extrapolation rendering consistency with extrapolation evaluation datasets generated by uniformly sampling positions and orientations in each scene from Replica. The quantitative resutls, comparing surface reconstruction, interpolation, and extrapolation rendering, are shown in Tab. I, our method shows the competitive surface reconstruction results with the state of the art, and the best performance in terms of interpolation rendering but fails to compete with the extrapolation rendering results of the NeRF-based M2Mapping. We validate our method is capable of capturing more precise geometric details in slim objects, as shown in Fig. 4. In the extrapolation views shown in Fig. 5, our method can retain more high-frequency textures (sharper spots) while NeRF-based methods show more natural transition (smoother lighting).
IV-C2 FAST-LIVO2 Datasets
For real-world scenes, we evaluate generalized types of trajectory datasets collected with a camera and a LiDAR and use the localization results from FAST-LIVO2 [5] as the input poses, as shown in the top row of Fig. 1 and Fig. 6. We adopt a train-test split for interpolation rendering evaluation, where every 8th photo is used for the test sets and the rest for the train sets [8], and the quantitative results are shown in Tab. II. Our method yields the best rendering performance overall, and also returns high-granularity surface reconstructions and considerable extrapolation rendering results across generalized scenes, as shown in Fig. 1 and Fig. 6.
TABLE II: Quantitative results on the FAST-LIVO2 datasets.
Metrics Methods Campus Sculpture Culture Drive Station SYSU CBD Avg.
[…]
3DGS 0.849 0.769 0.726 0.780 0.853 0.789 0.825 0.799
2DGS 0.839 0.730 0.611 0.741 0.798 0.808 0.760 0.755
PGSR 0.836 0.745 0.595 0.762 0.828 0.806 0.735 0.758
M2Mapping 0.834 0.729 0.727 0.764 0.809 0.789 0.850 0.786
Ours 0.858 0.774 0.797 0.788 0.861 0.841 0.892 0.830
[…]
3DGS 31.310 24.128 21.764 25.837 26.859 28.798 22.078 25.825
2DGS 30.611 22.654 19.218 24.847 24.756 27.517 22.022 24.518
PGSR 30.648 23.302 18.438 25.382 25.504 27.563 20.662 24.500
M2Mapping 30.681 23.453 24.695 25.941 25.859 28.011 26.655 26.471
Ours 31.697 24.196 25.074 25.657 27.065 28.532 26.696 26.988
We compare the qualitativer rendering results in a free trajectory (Drive), as shown in Fig. 8, where our method outperforms the others in retaining detailed textures of the floor with the help of geometric regularization. It is noted that both the LiDAR-augmented NeRF-based method (M2Mapping) and 3DGS-based method (Ours) give physically grounded rendering results, while the NeRF-based method shows more complete rendering at the view lacking observations and 3DGS-based method shows more detailed rendering overall. We also compared the reconstruction results over the diverse types of methods, as shown in Fig. 7. It is shown that LiDAR-based methods (VDBFusion, SHINE-Mapping, M2Mapping, Ours) offer far more detailed geometry than the vision-only methods (2DGS, PGSR).
(a) VDBFusion
(b) SHINE-Mapping
(c) 2DGS
(d) PGSR
(e) M2Mapping
(f) Ours
Figure 7: The surface reconstructions in the FAST-LIVO2 Campus dataset.
(a) InstantNGP
(b) 3DGS
(c) 2DGS
(d) PGSR
(e) M2Mapping
(f) Ground Truth
(g) RR
(h) RR+CR
(i) Ours (RR+SR)
Figure 8: The qualitative comparative results in the FAST-LIVO2 Drive dataset. We compare different geometric regularization strategies represented in Rendering Regularization (RR), Center Regularization (CR), and Shape Regularization (SR).
(a) Random Initialization (0 iteration)
(b) Random Initialization (30000 iteration)
(c) SDF-aided Initialization (0 iteration)
(d) SDF-aided Initialization (30000 iteration)
(e) SDF-aided Initialization with Shape Reg. (30000 iteration)
Figure 9: The ablation study of the SDF-aided Gaussian initialization and shape regularization in the FAST-LIVO2 Station scene. From left to right are the rendered color, depth, and normal images.
(a) 0 iter.
(b) 2000 iter.
(c) 30000 iter.
Figure 10: The ablation study of the color initialization in the Replica Office0 dataset. The first row shows the results without color initialization, and the second row shows the results with color initialization.
IV-D Ablation Study
IV-D1 Initialization
In this section, we only apply the render regularization (Eq. 5) as a pure image supervision to address the importance of the proposed Gaussian initializations (Sec. III-C) in converging a good structure of Gaussians. As shown in Fig. 9, the original initialilzation, with random sampling position from the input point clouds and random attributes initialization, falls into unreasonable structures (Fig. 9(b)). While, the proposed SDF-aided physically-grounded initialization gives a structure that more in line with the real scenarios. A color initialization assigns a proper appearance to the Gaussians, which is essential to stabilize the structure of Gaussians. As shown in Fig. 10, the inited Gaussian splats with proper structure but wrong color are deviated during early training.
IV-D2 Geometric regularization
We compare the geometric regularization methods between the render regularization [9], center regularization and shape regularization (Sec. III-D2), as shown in Fig. 8(g)-8(i). The render regularization (PSNR: 27.24, SSIM: 0.805, Fig. 8(g)) improves the geometrically consistency to a extent, but still produces false structure on the flat surfaces under the scenes with less multi-view constrains. The center regularization (PSNR: 28.60, SSIM: 0.861, Fig. 8(h)) shows its limited improvement in aligning Gaussians with the SDF field, still producing blur rendering on the ground. And our proposed shape regularization (PSNR: 28.66, SSIM: 0.866, Fig. 8(i)) shows better alignment with the surface to retain the details of the bricks. As shown in Fig. 9(e)’s rendered normal image, the shape regularization gives a more detailed and physically grounded structure of Gaussians for shaper rendering results.
IV-E Efficiency Analysis
The training time for the Replica Room-2 scene takes 16.6 minutes and the average rendering time for a 1200x680 image takes about 9.9 milliseconds and reaches 101.1 frames per second (FPS). We evaluate the efficiency of the compared NeRF-based method, M2Mapping, which takes 53.9 milliseconds for rendering a 1200x680 image (18.4 FPS). As our method trains both the NSDF and 3DGS, it takes more training time than the other 3DGS-based methods, like 2DGS (9.1 minutes / 103.2 FPS) but the inference time is comparable to the other methods and shows a significant improvement over the NeRF-based methods.
V Conclusion
In this paper, we addressed the geometric inconsistency challenges of Gaussian splatting in robotics applications by proposing a unified LiDAR-visual system that synergizes Gaussian splatting with neural signed distance fields. We leverage the NSDF to provide a physically grounded Gaussian initialization and effective shape regularization for geometrically consistent rendering and reconstruction. Extensive experiments demonstrate our method’s superior reconstruction accuracy and rendering quality across diverse trajectories. However, comparative analysis reveals the limitations in extrapolative novel view synthesis capabilities contrasted with NeRF-based frameworks. Therefore, we aim to tackle this limitation in future work by exploring advanced neural rendering techniques.